a parallel plate capacitor is given a charge of 3 micro columb. a sheet ofdielectric of dielectric constant 3, is inserted so that it completely fills the gap between theplates. the induced charge in micro coulomb on each face of the dielectric sheet is numerically

Grade:12

1 Answers

arpit juneja
6 years ago
Qind=Qext(1-1/k) where k is dielectric constant, Qind is induced charge on on dielectric sheet, Qext is charge on capacitor after inserting dielectric
Qext = k (CV) =3*3=9
Qind = 9(1-1/3)=6 microcoulomb